OpenStudy (anonymous):

Solve graphically x = 4 and 3x - 2y = 10. Write down the co-ordinates of the point of intersection of the two lines.

OpenStudy (anonymous):

12 - 2y = 10 2y = 2 y=1 ( 4 , 1 )

OpenStudy (anonymous):

3x-2y=10, 3(4)-2y=10 12-2y=10 12-10=2y 2=2y y=1 therefore the co ordinates are (4,1).